OpenStudy (anonymous):

Solve the system of linear equations below x+y=4 2x+3y=0

OpenStudy (matt101):

Rearrange the first equation to get y=4-x. Now we can sub this value of y into the second equation and solve for x: 2x+3(4-x)=0 2x+12-3x=0 -x=-12 x=12 Now we can plug x=12 into either of our original equations to find y. I'll use the first equation: 12+y=4 y=-8 And that's it!

OpenStudy (arabpride):

put into y = mx + b form x+y=4 -x -x y = -x + 4 2x+3y=0 -2x -2x 3y = -2x + 0 /3 /3 y = -2/3 +0 y = -2/3

OpenStudy (anonymous):

Thank you!

OpenStudy (arabpride):

Welcome *^.^*